Circular No. 5756

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 J. C. Wheeler and A. Clocchiatti, University of Texas, report: “Spec-

trograms (range 500-850 nm; dispersion 0.7 nm/pixel) obtained with the Imaging Grism Instrument on the 0.76-m telescope of McDonald Observatory on Apr. 6.3 and 7.3 UT show the onset of strong Hα and distinct P-Cyg profiles. The raw supernova spectrum was crudely corrected for instrumental effects by observing a nearby DA dwarf (GRW+705824 = EG102), fitting a polynomial to its continuum, and dividing the raw supernova spectrum by that polynomial. The result reveals that the spectrum on Apr. 7.3 has an emission peak at 654.5 nm of 29 nm FWHM with a broad minimum at 616.3 nm. The emission is roughly six times stronger than the absorption, appears to have sharp blue and red edges at 637 and 681 nm, respectively, and is asymmetric with a stronger red wing. This emission is most readily identified with Hα. The lack of deep absorption may indicate fluorescence from circumstellar material, which may affect the location of the minimum. Another feature shows a P-Cyg profile with a maximum at 588.4 nm and a well-defined minimum at 566.5 nm; this feature could be either He I 587.6 nm or Na D. For He I, the minimum represents a velocity of 10 760 km/s. The emission is approximately as strong as the absorption. The peak also shows a narrow absorption line of interstellar Na D. Another broad mimimum appears at ∼ 500 nm. Pending more accurate reduction, the spectrum resembles that of the type-II linear supernova 1979C on 1979 May 28 (Branch et al. 1981, Ap.J. 244, 780), although identification of the nature of the event at this time is premature.”

COMET MUELLER (1993a) Total magnitude estimates, visual except those by Mikuž (cf. IAUC

5699): Feb. 1.17 UT, 13.3 (H. Mikuž, Ljubljana, Slovenia, 0.20-m f/2 Baker-Schmidt camera + V filter + ST-6 CCD); 14.58, 13.3 (A. Nakamura, Kuma, Japan, 0.60-m reflector); 18.36, 13.2 (A. Hale, Las Cruces, NM, 0.41-m reflector); Mar. 10.81, 13.6 (Mikuž); 15.90, 12.7 (T. Vanmunster, Landen, Belgium, 0.35-m reflector); 26.93, 13.5 (Mikuž).

U GEMINORUM Corrigendum. On IAUC 5722, line 6, for 19.54, 11.1 read 19.34, 11.1

1993 April 8 Daniel W. E. Green